If a rocket is fired at an initial velocity of 350 meters per second, what is the maximum height it will atain and after how many seconds will it hit the ground? My teacher left me with this equation: h= ho+vot-1/2gt^2 where h= ending height, ho= starting height, vo= initial velocity, and g is gravity (9.8 meters per sec).
Answer by stanbon(75887)

If a rocket is fired at an initial velocity of 350 meters per second, what is the maximum height it will atain and after how many seconds will it hit the ground? My teacher left me with this equation: h= ho+vot-1/2gt^2 where h= ending height, ho= starting height, vo= initial velocity, and g is gravity (9.8 meters per sec).
---------
h(t)= -(1/2)9.8t^2+350t+0
When the rocket is back on the ground its height will be zero.
-4.9t^2+350t=0
-t(4.9t-350)=0
t=0 or 4.9t=350
t=0 or t=71.43 seconds
So the rocket was on the ground at time t=0
and the rocket will be on the ground after 71.43 seconds
----------------------
Cheers,
Stan H.